OpenStudy (anonymous):

Ned currently has a balance of $1,006.13 in an account he has held for 23 years. He opened the account with an initial deposit of $612. What is the simple interest rate on the account?

OpenStudy (anonymous):

i really need help finding this answer.. :[ it's the only way i can graduate..

OpenStudy (anonymous):

1006.13 - 612 = 394.13 (this is your interest) The formula you use is I = PRT, but you are looking for the rate, so change the formula to read R = I/PT. This is what we know... I = 394.13 P = 612 T = 23 Lets sub it in... R = I/PT R = 394.13 / (612)(23) R = 394.13 / 14076 R = .0279 rounded = .028 Now you have to turn the rate into a percent... .028 x 100 = 2.8% check.... I = PRT 394.13 = (612)(.028)(23) 394.13 = 394.128 rounded = 394.13 (correct) Ned has an interest rate of 2.8% Do you have any questions ?

OpenStudy (anonymous):

no thank you so very much:D

OpenStudy (anonymous):

your welcome :)